a father is three times as old as his son .5years ago fathers age was four times his son age then.find their present ages.
Answer by balthazar612(14)

(x + 4) = 3(y + 4)
(x - 6) = 7(y - 6)
---
x + 4 = 3y + 12
x - 6 = 7y - 42 ---
put the system of linear equations into standard form
---
x - 3y = 8
x - 7y = -36
Subtracting equati0n 2 from 1
4y = 44 Y = 11
Putting value of y in equation 1 we
get
X-3(11) =8
X -33 =8
X= 8+33 X= 41
answer:
x= father = 41
y= son = 11